OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#6778 — bea13 cor13 dra13 fer13 gar13 jol13 men13 pra13 smg13 sml13 stm13
Scheduled Maintenance Report for Web Cloud
Completed
We have a problem on the communication between the net center and the first NRA.

We are looking for the origin of the problem.

Update(s):

Date: 2012-05-25 10:43:56 UTC

Our tube was completely weldd again and we applyed some
bitumen on it. Everything is working again.


We had an appointment yesterday with France Telecom for
an alignment problem on the securisation fiber, but
the appointment was reported to Thursday May 31st.


So the next week, we are going to set up the securisation
on the second Marseille loop. This kind of problem will
be then invisible.

We are sorry for the non availability time.

Date: 2012-05-25 10:37:07 UTC
The links have just remounted.

The NAR are pinging again.

We are waiting for the end in order to close the incident
and have more information after that.

Date: 2012-05-25 10:35:59 UTC
The weldings are in progress. Some fibers are already UP.
We still have 144 to weld.
The reestablishment deadline for everybody
is expected to 12 hours.

Date: 2012-05-25 05:16:33 UTC
In the end, the temporary sealing has not been performed ...

The truck has been moved, access to the connexion box is available again.


Date: 2012-05-25 02:11:06 UTC
The trench was re-opened.It's not possible to get slack on the fiber because a truck is blocking the access ...

A temporary reparation will begin. There are 144 fibers to be welded. For the moment the teams are preparing the \"field\".

Date: 2012-05-25 00:26:00 UTC
Well ... re-routing is not possible.
The path is being re-opened to retrieve the scabbard abyss and start welding.

Date: 2012-05-25 00:24:33 UTC
The cut-off has been localized.Civil engineering works has cut a sleeve of 10m long. Bitumen was poured on top. The reparation is not possible.
The network teams are looking in order to find free fibers in the neighbor sleeves to make a temporary re-routing.

Date: 2012-05-25 00:16:02 UTC
The cut-off was \"prelocalised\" 800m from netcenter.
The teams are on the spot and looking for the exact location.

So for the moment we do not know the type of the cut-off and how long the reparation will take.

Date: 2012-05-24 20:44:57 UTC
The interruption was localized. The teams are on the spot.

Date: 2012-05-24 19:56:38 UTC
The first tests show a problem on the fiber
section. The technicians will be on the spot
at 20:30 and we will have more informations
about the interruption.

Date: 2012-05-24 19:53:10 UTC
The tickets were opened with different operators.
Posted May 24, 2012 - 19:52 UTC